OpenStudy (anonymous):

Will medal The sum of the ages of Andres and Riley is 80 years. 10 years ago, Andres's age was 2 times Riley's age. How old is Andres now?

OpenStudy (anonymous):

50. 50 + 30 = 80 50-10 = 40 30 - 10 = 20 40 / 2 = 20

OpenStudy (anonymous):

How did you get 50?

OpenStudy (anonymous):

I did simple math. 40 + 40= 80 50 + 30 = 80 60 + 20 = 80 70 + 10 = 80 Then I figured what would fit the rest of the problem

OpenStudy (anonymous):

So Andres is 50

OpenStudy (anonymous):

Yes!